How many gallons of a 10% antifreeze solution and a 25 % Antifreeze solution must be mixed to make 3 gallons of a 20 % antifreeze solution? Solve with equations
Answer by tommyt3rd(5050)

How many gallons of a 10% antifreeze solution: 0.10x
and a 25 % Antifreeze solution must be mixed: 0.25y
to make 3 gallons of a 20 % antifreeze solution: (3)(0.20)=0.6

but x+y=3 (so y=3-x)
and we write


0.1x+0.25(3-x)=0.6
which gives us x=1 gallon



:)